OpenStudy (anonymous):

Simplify square root of 54 a. square root of 18 b. 6* square root of 3 c. 3* square root of 6 d, 9* square root of 6

OpenStudy (anonymous):

here we go again, \[\sqrt{54}=\sqrt{9}*\sqrt{6}=3\sqrt{6}\]